A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1900–1909

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.01

Title word cross-reference

1870 [Ano05a]. 1905 [Smi05b]. 1906 [Tit06].

21st [Ano08e].

Absence [Ano08f]. Accompanied [Din01]. Acipenser [Sto00]. Address [Ano00a, Ano08a]. Adult [Fie07]. Advances [Mea03]. Aerating [Ano06f]. After [Den07]. Aid [Ful06, O’M05]. Air [Mar03]. amended [Ano00c, Ano01b, Ano02b, Ano03b, Ano04b, Ano05c, Ano06c, Ano07c]. America [Col05, Tit04, Tit05a]. American [Ano00b, Ano01a, Ano02a, Ano03a, Ano04a, Ano05b, Ano06b, Ano07b, Ano08e, Ano08b, Ano05a, Ano06a, Ano07a, Ano08c, Lea06]. among [Mar03]. Angling [Bar03]. Annual [Ano08a]. Apparatus [Ano05e]. Aquatic [Gre04, Par01, Pon02]. Argyrosomus [Wag08]. Arlington [Ano08e].

Arrested [Str00]. Artificial [Ber08, Cla02]. Assist [Eva07]. Associations [Eva07]. Assorting [Str06]. Away [Bir07a].

Basin [Bea08, Sto01]. Bass [Bar08, Bee06, Bir07a, Cla05, Dea02, Fea03, Lam04, Lam00, Lam02, Lea03, Lov04, Lyd02, Lyd04, Lyd06a, Rip08, Str03, Str06, Str08, Tal05, Wor04]. Bay [Ano03d]. Bear [Hub03]. Beds [Ano03d]. before [Smi06]. Best [Bar03, Lyd06b]. Biology [Pon02]. Black [Bar08, Bee06, Bir07a, Lam00, Lam02, Lea03, Lov04, Lyd02, Rip08, Str06, Str08]. Blood [Hen03]. Box [Lea06]. Breeding [Mea00]. Brood [Str06]. Brook [Ano00d, Hub03, Mar01, Mar02, Tho00, Tho01, Whi05, Woo00]. Bryant [Pea04]. Bureau [Tit05b, Tit06].

C [Ano08e]. Cannibalism [Str06]. Cans [Mar04]. Capacity [Hub03]. Care [Ano00d, Bul05, Joh06]. Carp [Bar00, Bar01, Bar03, Bar05, Bar08, Col05]. Ceding [Jos05]. Cement [Mar02]. Champlain [Sto01]. Charr [Joh04]. Clam [Mea00]. Close [Jam00]. Cold [Mar04]. Collecting [Dow05]. Commercial [Cla00, Whi03, Wor04]. Commission [Whi03, Mar04]. Commissioners [Mea05]. Committee [Ano08e, Ano00g, Ano01f, Ano02e, Ano03f, Ano04e, Ano05h, Ano06e, Ano07e, Ano08g, Atk06b]. Common [Gil08]. Company [Ano03d]. Complete [Ano08c]. Concerning [Coh01]. Congress [Smi05b, Smi06]. Connection [Cla00, Lyd04]. Constitution [Ano00c, Ano01b, Ano02b, Ano03b, Ano04b, Ano05c, Ano06c, Ano07c, Ano08d]. Containing [Mar03]. Contrivances [Woo00]. Control [Jos05]. Controlling [Par01]. Cooking [Bar03]. Corstruction [Lea04]. County [Ano03d]. Creek [Lyd04, Lyd06a, Eve05]. Cultivation [Smi05a, Tow07]. Cultural [Lea04]. Culture [Ano05d, Bea07, Car04, Cla05, Dea02, Gre04, Hen00, Hen01, Lyd02, Mea01, Mea03, Mea05, Mee05, Moo04, O’M05, Smi07, Str02, Str03, Tit05b, Tit06]. Culturist [Eva07].

D [Ano08e]. Dagget [Ano03d]. Danger [Mar04]. date [Ano00c, Ano01b, Ano02b, Ano03b, Ano04b, Ano05c, Ano06c, Ano07c]. December [Ano05a, Ano06a, Ano07a]. Delaware [Mee07]. Department [Bea04]. Description [Lea08, Woo00]. Difficulties [Bea07]. Discouragements [Dea02]. Discussion [Ano00d, Ano05d, Bar01, Lam02, Woo00]. Disease [Mar01, Mar02]. Diseases [Atk01]. Display [Ano05e]. Dissolved [Bir06, Mar03]. Distributed [Mee06a]. Distribution [Dow05]. Done [Joh04]. Dressed [Rei06]. During [Joh04, Tit05b, Tit06].

Early [Atk05, Fea03]. East [Ano00d]. Eastern [Joh04]. Economic [Wag08]. Eel [Gil08]. Effect [Tho02]. Egg [Tho02]. Eggs [All05, Atk07, Bir07a, Bow05, Dow05, Hub03]. Eighth [Ano08f]. etc [Sea02].

European [Smi07]. Excess [Mar03]. Executive [Ano00g, Ano01f, Ano02e, Ano03f, Ano04e, Ano05h, Ano06e, Ano07e, Ano08g]. Exhibition [Woo00]. Experimental [Mea05]. Experiments [Atk06a, Ber08, Bir07a, Hen04, Mea01, Moo04, Rip08, Ros06, Tit05b, Tit06, Mar04]. Exposition [Bea04]. Expression [Ber08].

Factor [Par01, Woo01]. Failure [Den08]. Farm [Lov03, Str02]. Fasting [Atk06a]. Fatality [Mar03]. Feeding [All05, Ano00d, Atk05, Hen04, Joh06, Tho02]. Fertilization [Ber08, Str00]. Few [Lam02]. Fifty [Smi06]. Fifty-Ninth [Smi06]. First [Smi06]. Fiscal [Tit06]. Fish [Ano00a, Ano08e, Atk01, Bea07, Ber08, Bre00, Bry01, Coh01, Den05, Den08, Eva07, Hen01, Hen03, Hen05, Jam00, Jam01, Lea07, Lov03, Mee06a, O’M05, Smi07, Str02, Tit05b, Tit06, Wag08, War03, Whi03, Bea04, Bre00]. Fisheries [Ano00b, Ano01a, Ano02a, Ano03a, Ano04a, Ano05a, Ano05b, Ano06a, Ano06b, Ano07a, Ano07b, Ano08c, Ano08e, Ano08b, Bry01, Cok07, Eva07, Fie07, Lea06, Mee04, Ols07, Smi05b, Smi07, Tit05a, Mea05, Tit05b, Tit06]. Fisherman [Rot06]. Fishery [Cob08, Smi06, Tit04]. Fishes [Atk04, Cla00, Den07, Hen02, Mar03, Tow07]. Fishing [Din05, Smi04]. Fishway [Buc06]. Food [Atk03, Bar00, Bar08, Den07, Hen02, Hen03, Jam01, War03]. Foreign [Atk06b]. Forest [Mar04, Whi03]. Free [Ano00d]. Free-Town [Ano00d]. Fresh [Pon02]. Friend [Bar05]. Frog [Mee05]. Fry [Atk05, Atk06a, Hen04, Joh06, Lea07, Tho00, Woo01]. Furnishing [Bar08].

Game [Ano00a, Bea04, Bre00, Eva07, Hen02, Mar04, Whi03]. Gases [Bir06]. Gold [Ber08]. Golden [Eve05, Tho03]. Grand [Buc06]. Grant [Ano00a]. Grayling [Bab00, Hen00]. Great [Dow05, Cla00, Eva07, Jos05, Rei06]. Green [Hub03]. Growth [Mea00, Tho02]. Guano [Cok07].

Habits [Lam00, Lyd02, Mea00, Sto00, Str08]. Harbor [Mar04]. Hatcheries [Woo00]. Hatchery [Mar04, Ano00d]. Hatching [Bir07a, Cla07, Dow05, Ful08, Mee06b, Sto01, Wor04]. Hawaiian [Cob08]. Herring [Rei06]. Hints [Bar03, Hen00, Hen01]. History [Gil08]. Hon. [Pea04]. Hotel [Ano08e]. House [Ful08]. Huntoon [Ano03d].

Identification [Rei06]. Illinois [Bar00, Coh01]. Illustrated [Din05]. Importance [Wag08]. Inbreeding [Syk02]. Increase [Gre08]. Index [Ano00e, Ano01c, Ano02c, Ano03c, Ano04c, Ano05f, Gun01]. Industry [Cok07]. Influence [Eva07]. Inland [Bir07b, Hen05, Mea05]. Interest [Mee04]. International [Eva07, Ols07, Smi05b]. Island [Mea05].

January [Mee06a]. Japan [Smi04]. Jar [Cla07]. Jars [Mee06b]. July [Mee06a].

Lack [Str00]. Lake [Bir07b, Gre08, Sto00, Buc06, Gre08, Sto01]. Lakes [Bir06, War03, Cla00, Eva07, Jos05, Rei06]. Lantern [Din01, Din05]. Large [Lam00, Lea03, Str08]. Large-Mouth [Lam00, Str08, Lea03]. Larger [Pon02]. Latter [Eva07]. Laws [Coh01, Den07, Jam01]. Legal [Rei06]. Legislation [Den08, Smi06]. Lessons [Mee07]. Letter [Ano03d, Bee06]. Life [Gil08, Joh06, Par01]. List [Ano00f, Ano01d, Ano02d, Ano03e, Ano04d, Ano05g, Ano06d, Ano07d, Ano08c]. Live [Atk03]. Live-Food [Atk03]. Lobster [Fie07, Mea01, Mea03, Mea05]. Loss [Dow05]. Louis [Bea04].

M [Ano00a]. made [Mar04]. Main [Str03]. Maine [Din01]. Maintain [Fie07]. Making [Den07]. Man [Par01]. Manipulation [Atk07]. Marcos [Lea03, Lea08]. May [Eva07]. Meeting [Ano08e, Ano08f]. Members [Ano00f, Ano01d, Ano02d, Ano03e, Ano04d, Ano05g, Ano06d, Ano07d]. Membership [Ano08c]. Methods [Cla00, Lea04, Lea08, Tho00]. Mich. [Ano00a]. Michigan [Cla05, Bab00]. Migration [War08]. Mill [Lyd04, Lyd06a]. Mode [Bar03]. Modern [Ful08]. Morse [Ano00a]. Mountain [Hen02]. Mouth [Cla05, Lam00, Rip08, Str08, Lea03]. Mr. [Ano00a, Ano05d]. Muskalonge [Bea08]. Muskellunge [Nev01]. Mutilated [Rei06].

Nation [Smi04]. National [Din05, Jos05]. Native [Whi05]. Necessary [Str03]. Necessity [Den07, Fie07]. Need [Ols07]. Neglected [Atk04]. Nests [Bir07a]. Ninth [Smi06]. Northville [Cla05]. Note [Ano01e, Ano08f]. Notes [All05, Bow05, Car04, Cla05, Cla07, Fea03, Lyd04, Mar04, Tho01, War03, Whi03].

Observations [RW07, Smi07]. Officers [Ano00g, Ano01f, Ano02e, Ano03f, Ano04e, Ano05h, Ano06e, Ano07e, Ano08g]. Ohio [Bea08]. Ontario [Gre08]. Order [Fie07]. Organized [Ano05a, Ano06a, Ano07a]. Original [Woo00]. other [Wor04]. Oyster [Ano03d]. Oysters [Ano03d].

p.m [Ano08e]. Pacific [Dow00, War08]. Paint [Lyd06b]. Paramount [Smi04]. Parasites [War08]. Parent [All05]. Parents [Bir07a]. Park [Din05]. Passing [Whi05]. past [Joh04, Tit05b]. Peculiarities [Str08]. Pennsylvania [Jam01, Mee04, Mee06a]. Perch [Bul05, Dow05]. Peru [Cok07]. Pike [Dow05]. Pike-Perch [Dow05]. Plan [Lam04]. Planting [Lea07, Sea02]. Plants [Gre04, Pon02]. Points [Lam02, Str03, War08]. Policy [Jos05]. Politics [Eva07]. Pond [Ano05d, Gre04, Joh06, Lea04, Syk02]. Pond-Reared [Syk02]. Ponds [Lam04, Lea04, Mar02, Tow07]. Portland [Ano00a]. Possibilities [Wor04]. Potomac [Tal05]. Power [Bry01]. Practical [Bea07, Hen01, Str07]. Preface [Ano00h]. President [Ano08a]. Press [Ada00]. Prevent [Str06]. Problem [Atk03, Eva07, Mea05]. Produced [All05]. Product [Bar00]. Production [Tho02]. Progress [Moo04, Tit05b, Tit06]. Propagation [Bee06, Bre00, Bul05, Cla00, Cla02, Dow00, Dow04, Ful06, Lea03, Nev01, Ros06]. Proper [Eva07]. Propogation [Lea08]. Protection [Den05, Den07, Dow04, Fie07, Ful06, Hen05, Jam01]. Protective [Den08, Eva07]. Purposes [Rei06].

Quality [Woo01]. Question [Lea06]. Questions [Lea06]. Quinnat [Bow05].

Rainbow [RW07, Ros06, Sea02]. Raising [Bir07a, Joh04]. Re [Ano08f]. Reared [Syk02]. Rearing [Rip08, Woo01]. Reasons [Den08]. Reference [All05]. Reforestation [Rot06]. regarding [Ano03d, RW07]. Region [Hen02]. Regulate [Bry01]. Regulation [Eva07]. Relation [Bre00]. Relations [Atk06b]. Relative [Hub03]. Remarks [Sea02, Smi05a]. Remedies [Den08]. Reminiscences [Tit04, Tit05a]. Report [Ano05d, Atk06b]. Resort [Din05]. Respiration [Bir07b]. Responding [Ada00]. Results [Cla00]. Resume [Joh04, Tho00]. Rhode [Mea05]. Richardson [Wag08]. Rights [Cob08]. River [Mee07]. Rocky [Hen02]. Role [Pon02]. Rough [Hub03]. rubicundus [Sto00].

Salmon [Atk07, Bow05, Dow00, Hub03, War08]. Salmonoid [Atk05]. Salt [O’M05]. Samish [Ano03d]. Samples [Ano03d]. San [Lea03, Lea08]. Screens [Lyd06b]. Season [Jam00]. Seed [Ano03d]. Seen [Bar05, Lov04]. Segmentation [Str00]. Select [Lov03]. September [Ano08e]. Session [Smi06]. Shad [Mee07]. Shipping [Mar04]. shown [War08]. Sierras [War03]. Size [Ano08e]. Slides [Din01, Din05]. Small [Cla05, Rip08, Tow07]. Small-Mouth [Rip08]. Society [Ano08e, Ols07, Ano05a, Ano06a, Ano07a, Ano08c, Lea06, Ano00b, Ano01a, Ano02a, Ano03a, Ano04a, Ano05b, Ano06b, Ano07b, Ano08b]. Solution [O’M05]. Some [Bar03, Bea07, Ber08, Den08, Dow04, Fea03, Lea08, Lyd04, Ros06, Sea02, Smi07, Str08, War03, War08, Whi03]. South [Lam00, Tit04, Tit05a]. Spawning [Lam00, RW07, Sto00, Str08]. Species [Hub03, Lov03, RW07, Wor04]. Specimens [Rei06]. Sponge [Moo04, Smi05a]. Sportsman [Din01]. Spring [Mar04]. St [Bea04]. State [Ano00a, Bry01, Den07, Jos05, Whi03]. States [Den07]. Station [Lea08, Cla05, Lea03, Lea08, Lyd04, Lyd06a]. Status [Col05]. Stocking [Den07]. Stream [Buc06]. Striped [Fea03, Wor04]. Study [Atk01, Cok07, Whi03]. Sturgeon [Car04, Sto00, Sto01]. Successful [Cla02, Str03]. Suggested [Ano00d]. Suggestive [Den08]. Supervision [Jos05].

Taken [Ano03d]. Taking [Bir07a, Bow05, Bry01]. Tests [Str07]. Theory [Str07]. Thirty [Ano08f]. Thirty-Eighth [Ano08f]. Thoughts [Dow04]. Three [Str03]. Thymallus [Bab00]. Time [RW07, Sea02]. Titcomb [Ano05d]. Town [Ano00d]. Transactions [Ano08f, Ano00b, Ano01a, Ano02a, Ano03a, Ano04a, Ano05b, Ano06b, Ano07b, Ano08b]. Transportation [Hub03]. Tribute [Pea04]. tricolor [Bab00]. Trip [Ano00d]. Troughs [Bir07a]. Trout [All05, Ano00d, Eve05, Gre08, Hub03, Mar01, Mar02, RW07, Ros06, Sea02, Syk02, Tho00, Tho01, Tho03, Whi05, Woo00, Woo01]. Tube [Ano06f]. Tullibee [Wag08, Wag08]. two [Hub03]. Type [Mee06b].

United [Den07]. Universal [Bea04]. upon [Eva07]. Usage [Hub03]. Use [Woo00, Lea08]. Utilization [Atk04].

Value [Bar00, Bar08, Gre04]. Values [Whi03]. Varieties [RW07]. Vermont [Car04]. Vienna [Smi05b]. Virility [All05]. Volcano [Eve05]. vs [Lea07, Str00, Str07].

Warden [Ano00a, Bre00]. Washington [Ano08e, Ano03d]. Water [Ano06f, Mar03, Pon02, Woo01]. Waters [Bar00, Bir06, Den07, Hen05, Joh04]. Way [Bea07]. Western [Joh04]. Whitefish [Cla02, Dow04, Gre08, Rei06]. Wisconsin [Bir06, Nev01]. Work [Bre00, Eva07, Joh04, Lyd04, Mea05, Mee04, Mee07, Whi03].

Year [Cla02, Joh04, Mee04, Tit05b, Tit06]. Yellow [Bul05]. Yellowstone [Din05]. York [Mar04, Whi03].

References

Adams:1900:RP

[Ada00] Fred J. Adams. Responding to the press. Transactions of the American Fisheries Society, 29(1):180–183, January 1900. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1905:NFP

[All05] George R. Allen. Notes on the feeding of parent trout, with reference to virility of eggs produced. Transactions of the American Fisheries Society, 34(1):122–123, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1900:AMG

[Ano00a] Anonymous. Address of Mr. Grant M. Morse, State Game and Fish Warden, Portland, Mich. Transactions of the American Fisheries Society, 29(1):164–167, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1900:TAF

[Ano00b] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 29(1):5–46, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1900:CAD

[Ano00c] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 29(1):191–192, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1900:DCF

[Ano00d] Anonymous. Discussion on the care and feeding of brook trout, suggested by the trip to the East Free-Town Trout Hatchery. Transactions of the American Fisheries Society, 29(1):168–172, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1900:I

[Ano00e] Anonymous. Index. Transactions of the American Fisheries Society, 29(1):4, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1900:LM

[Ano00f] Anonymous. List of members. Transactions of the American Fisheries Society, 29(1):184–190, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1900:OEC

[Ano00g] Anonymous. Officers for 1900–1901 and Executive Committee. Transactions of the American Fisheries Society, 29(1):2, January 1900. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1900:P

[Ano00h] Anonymous. Preface. Transactions of the American Fisheries Society, 29(1):3, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1901:TAF

[Ano01a] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 30(1):5–26, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1901:CAD

[Ano01b] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 30(1):174–175, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1901:I

[Ano01c] Anonymous. Index. Transactions of the American Fisheries Society, 30(1):4, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1901:LM

[Ano01d] Anonymous. List of members. Transactions of the American Fisheries Society, 30(1):163–173, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1901:N

[Ano01e] Anonymous. Note. Transactions of the American Fisheries Society, 30(1):3, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1901:OEC

[Ano01f] Anonymous. Officers for 1901–1902 and Executive Committee. Transactions of the American Fisheries Society, 30(1):2, January 1901. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1902:TAF

[Ano02a] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 31(1):5–40, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1902:CAD

[Ano02b] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 31(1):169–171, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1902:I

[Ano02c] Anonymous. Index. Transactions of the American Fisheries Society, 31(1):3, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1902:LM

[Ano02d] Anonymous. List of members. Transactions of the American Fisheries Society, 31(1):157–168, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1902:OEC

[Ano02e] Anonymous. Officers for 1902–1903 and Executive Committee. Transactions of the American Fisheries Society, 31(1):2, January 1902. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1903:TAF

[Ano03a] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 32(1):5–42, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1903:CAD

[Ano03b] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 32(1):235–237, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1903:I

[Ano03c] Anonymous. Index. Transactions of the American Fisheries Society, 32(1):3, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1903:LHO

[Ano03d] Anonymous. Letter from huntoon oyster company regarding samples of seed oysters taken from oyster beds at Samish Bay, Dagget County, Washington. Transactions of the American Fisheries Society, 32(1): 109–111, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1903:LM

[Ano03e] Anonymous. List of members. Transactions of the American Fisheries Society, 32(1):221–233, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1903:OEC

[Ano03f] Anonymous. Officers for 1903–1904 and Executive Committee. Transactions of the American Fisheries Society, 32(1):2, January 1903. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1904:TAF

[Ano04a] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 33(1):5–42, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1904:CAD

[Ano04b] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 33(1):262–264, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1904:I

[Ano04c] Anonymous. Index. Transactions of the American Fisheries Society, 33(1):3, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1904:LM

[Ano04d] Anonymous. List of members. Transactions of the American Fisheries Society, 33(1):245–261, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1904:OEC

[Ano04e] Anonymous. Officers for 1904–1905 and Executive Committee. Transactions of the American Fisheries Society, 33(1):2, January 1904. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1905:AFS

[Ano05a] Anonymous. American Fisheries Society: Organized December, 1870. Transactions of the American Fisheries Society, 34(1):3, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1905:TAF

[Ano05b] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 34(1):7–46, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1905:CAD

[Ano05c] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 34(1):282–284, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1905:DMT

[Ano05d] Anonymous. Discussion on Mr. Titcomb’s report on pond culture. Transactions of the American Fisheries Society, 34(1):190–194, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1905:DA

[Ano05e] Anonymous. Display of apparatus. Transactions of the American Fisheries Society, 34(1):118–121, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1905:I

[Ano05f] Anonymous. Index. Transactions of the American Fisheries Society, 34(1):5, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1905:LM

[Ano05g] Anonymous. List of members. Transactions of the American Fisheries Society, 34(1):265–281, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1905:OEC

[Ano05h] Anonymous. Officers for 1905–1906 and Executive Committee. Transactions of the American Fisheries Society, 34(1):2, January 1905. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1906:AFS

[Ano06a] Anonymous. American Fisheries Society: Organized December, 1870. Transactions of the American Fisheries Society, 35(1):3, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1906:TAF

[Ano06b] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 35(1):7–43, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1906:CAD

[Ano06c] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 35(1):252–254, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1906:LM

[Ano06d] Anonymous. List of members. Transactions of the American Fisheries Society, 35(1):237–251, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1906:OEC

[Ano06e] Anonymous. Officers for 1906–1907 and Executive Committee. Transactions of the American Fisheries Society, 35(1):2, January 1906. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1906:TAW

[Ano06f] Anonymous. Tube for aerating water. Transactions of the American Fisheries Society, 35(1):188, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1907:AFS

[Ano07a] Anonymous. American Fisheries Society: Organized December, 1870. Transactions of the American Fisheries Society, 36(1):3, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1907:TAF

[Ano07b] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 36(1):7–100, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1907:CAD

[Ano07c] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 36(1):269–271, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1907:LM

[Ano07d] Anonymous. List of members. Transactions of the American Fisheries Society, 36(1):255–268, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1907:OEC

[Ano07e] Anonymous. Officers for 1907–1908 and Executive Committee. Transactions of the American Fisheries Society, 36(1):2, January 1907. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1908:AAP

[Ano08a] Anonymous. Annual address of the President. Transactions of the American Fisheries Society, 37(1):69–74, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1908:TAF

[Ano08b] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 37(1):9–67, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1908:CML

[Ano08c] Anonymous. Complete membership list: American Fisheries Society 1908. Transactions of the American Fisheries Society, 37(1):176–192, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1908:C

[Ano08d] Anonymous. Constitution. Transactions of the American Fisheries Society, 37(1):193–195, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1908:MCS

[Ano08e] Anonymous. Meeting of Committee on Size of Fish, American Fisheries Society: Washington, D. C, September 21st, 1908, 6 p.m., at Arlington Hotel. Transactions of the American Fisheries Society, 37(1): 164–175, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1908:NRA

[Ano08f] Anonymous. Note re the absence of the Transactions of the Thirty-Eighth Meeting. Transactions of the American Fisheries Society, 37 (1):196, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1908:OEC

[Ano08g] Anonymous. Officers for 1908–1909 and Executive Committee. Transactions of the American Fisheries Society, 37(1):2, January 1908. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1901:SFD

[Atk01] Chas G. Atkins. The study of fish diseases. Transactions of the American Fisheries Society, 30(1):82–89, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1903:LFP

[Atk03] Chas G. Atkins. The live-food problem. Transactions of the American Fisheries Society, 32(1):71–78, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1904:UNF

[Atk04] Charles G. Atkins. The utilization of neglected fishes. Transactions of the American Fisheries Society, 33(1):178–191, January 1904. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1905:EFS

[Atk05] Charles E. Atkins. The early feeding of salmonoid fry. Transactions of the American Fisheries Society, 34(1):75–89, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1906:EFF

[Atk06a] Charles G. Atkins. Experiments in fasting of fry. Transactions of the American Fisheries Society, 35(1):123–142, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1906:RCF

[Atk06b] Charles G. Atkins. Report of Committee on Foreign Relations. Transactions of the American Fisheries Society, 35(1):189–224, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1907:MSE

[Atk07] Charles G. Atkins. Manipulation of salmon eggs. Transactions of the American Fisheries Society, 36(1):218–222, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Babbitt:1900:MGT

[Bab00] A. C. Babbitt. Michigan grayling: (Thymallus tricolor). Transactions of the American Fisheries Society, 29(1):106–108, January 1900. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartlett:1900:VCF

[Bar00] S. P. Bartlett. The value of the carp as a food product of Illinois waters. Transactions of the American Fisheries Society, 29(1):80–87, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartlett:1901:DC

[Bar01] S. P. Bartlett. Discussion on carp. Transactions of the American Fisheries Society, 30(1):114–132, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartlett:1903:ACS

[Bar03] S. P. Bartlett. Angling for carp, and some hints as to best mode of cooking. Transactions of the American Fisheries Society, 32(1):47–57, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartlett:1905:CSF

[Bar05] S. P. Bartlett. Carp, as seen by a friend. Transactions of the American Fisheries Society, 34(1):207–216, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartlett:1908:VCF

[Bar08] S. P. Bartlett. Value of carp as furnishing food for black bass. Transactions of the American Fisheries Society, 37(1):85–91, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1904:FGD

[Bea04] Tarleton H. Bean. Fish and Game Department of the Universal Exposition at St. Louis. Transactions of the American Fisheries Society, 33 (1):55–59, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1907:SPD

[Bea07] Tarleton H. Bean. Some practical difficulties in the way of fish culture. Transactions of the American Fisheries Society, 36(1):184–192, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1908:MOB

[Bea08] Tarleton H. Bean. The muskalonge of the Ohio Basin. Transactions of the American Fisheries Society, 37(1):145–151, January 1908. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beeman:1906:LPB

[Bee06] Henry W. Beeman. Letter on the propagation of black bass. Transactions of the American Fisheries Society, 35(1):182, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berkhouse:1908:SEA

[Ber08] Jerry R. Berkhouse. Some experiments on the artificial expression and fertilization of gold fish. Transactions of the American Fisheries Society, 37(1):142–144, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Birge:1906:GDW

[Bir06] E. A. Birge. Gases dissolved in the waters of Wisconsin lakes. Transactions of the American Fisheries Society, 35(1):143–163, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Birge:1907:ERB

[Bir07a] Edward A. Birge. Experiments in raising black bass by taking the nests away from the parents and hatching the eggs in troughs. Transactions of the American Fisheries Society, 36(1):154–158, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Birge:1907:RIL

[Bir07b] Edward A. Birge. The respiration of an inland lake. Transactions of the American Fisheries Society, 36(1):223–245, January 1907. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:1905:NTQ

[Bow05] Ward T. Bower. Notes on the taking of quinnat salmon eggs. Transactions of the American Fisheries Society, 34(1):232–238, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brewster:1900:RFG

[Bre00] C. E. Brewster. The relation of the Fish and Game Warden in the work of fish propagation. Transactions of the American Fisheries Society, 29(1):69–79, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bryant:1901:PSR

[Bry01] Edwin E. Bryant. The power of the state to regulate fisheries, and the taking of fish. Transactions of the American Fisheries Society, 30 (1):31–47, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buck:1906:FGL

[Buc06] W. O. Buck. The fishway at Grand Lake stream. Transactions of the American Fisheries Society, 35(1):233–236, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buller:1905:PCY

[Bul05] Nathan R. Buller. Propagation and care of yellow perch. Transactions of the American Fisheries Society, 34(1):223–231, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carter:1904:NSC

[Car04] E. N. Carter. Notes on sturgeon culture in Vermont. Transactions of the American Fisheries Society, 33(1):60–75, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1900:MRC

[Cla00] Frank N. Clark. Methods and results in connection with the propagation of commercial fishes for the Great Lakes. Transactions of the American Fisheries Society, 29(1):88–98, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1902:SYA

[Cla02] Frank N. Clark. A successful year in the artificial propagation of the whitefish. Transactions of the American Fisheries Society, 31(1): 97–106, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1905:NSM

[Cla05] Frank N. Clark. Notes on small mouth bass culture at the Northville, Michigan, Station. Transactions of the American Fisheries Society, 34(1):174–189, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1907:NNH

[Cla07] Frank N. Clark. Notes on a new hatching jar. Transactions of the American Fisheries Society, 36(1):159–169, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1908:HFR

[Cob08] John N. Cobb. Hawaiian fishery rights. Transactions of the American Fisheries Society, 37(1):160–163, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cohen:1901:CFL

[Coh01] Nat H. Cohen. Concerning fish laws in Illinois. Transactions of the American Fisheries Society, 30(1):133–136, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coker:1907:SGI

[Cok07] Robert E. Coker. A study of the guano industry and fisheries of Peru. Transactions of the American Fisheries Society, 36(1):248–251, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cole:1905:SCA

[Col05] Leon J. Cole. The status of the carp in America. Transactions of the American Fisheries Society, 34(1):201–206, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dean:1902:DBC

[Dea02] H. D. Dean. Discouragements in bass culture. Transactions of the American Fisheries Society, 31(1):153–155, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dennis:1905:FP

[Den05] Oregon Milton Dennis. Fish protection. Transactions of the American Fisheries Society, 34(1):133–138, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dennis:1907:NSM

[Den07] Oregon Milton Dennis. The necessity of the State making laws for the protection of food fishes after stocking waters by the State or United States. Transactions of the American Fisheries Society, 36(1): 103–104, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dennis:1908:SRF

[Den08] Oregon Milton Dennis. Some reasons for failure of fish protective legislation and some suggestive remedies. Transactions of the American Fisheries Society, 37(1):152–156, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dinsmore:1901:MSA

[Din01] A. H. Dinsmore. Maine and the sportsman: (accompanied by lantern slides). Transactions of the American Fisheries Society, 30(1):159– 162, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dinsmore:1905:YPN

[Din05] A. H. Dinsmore. Yellowstone Park as a national fishing resort: (illustrated by lantern slides). Transactions of the American Fisheries Society, 34(1):195–198, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Downing:1900:PPS

[Dow00] S. W. Downing. Propagation of the Pacific salmon. Transactions of the American Fisheries Society, 29(1):154–163, January 1900. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Downing:1904:WST

[Dow04] S. W. Downing. The whitefish; some thoughts of its propagation and protection. Transactions of the American Fisheries Society, 33(1): 104–110, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Downing:1905:CHD

[Dow05] S. W. Downing. Collecting, hatching and distribution of pike-perch: Why the great loss of eggs. Transactions of the American Fisheries Society, 34(1):239–255, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:1907:IPU

[Eva07] A. Kelly Evans. The influence of politics upon the work of the fish culturist, and how fish and game protective associations may assist the latter and the international problem of the proper regulation of the fisheries of the Great Lakes. Transactions of the American Fisheries Society, 36(1):207–217, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evermann:1905:GTV

[Eve05] Barton W. Evermann. The golden trout of Volcano Creek. Transactions of the American Fisheries Society, 34(1):148–155, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fearing:1903:SEN

[Fea03] D. B. Fearing. Some early notes on striped bass. Transactions of the American Fisheries Society, 32(1):90–108, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Field:1907:NPA

[Fie07] G. W. Field. The necessity of the protection of the adult lobster, in order to maintain the lobster fisheries. Transactions of the American Fisheries Society, 36(1):193–206, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fullerton:1906:PAP

[Ful06] Samuel F. Fullerton. Protection as an aid to propagation. Transactions of the American Fisheries Society, 35(1):59–86, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fullerton:1908:MHH

[Ful08] S. F. Fullerton. The modern hatching house. Transactions of the American Fisheries Society, 37(1):132–141, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gill:1908:LHC

[Gil08] Theodore Gill. Life history of the common eel. Transactions of the American Fisheries Society, 37(1):115–121, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1904:VAP

[Gre04] C. K. Green. Value of aquatic plants in pond culture. Transactions of the American Fisheries Society, 33(1):173–177, January 1904. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1908:IWL

[Gre08] Chester K. Green. The increase of whitefish and lake trout in Lake Ontario. Transactions of the American Fisheries Society, 37(1):82–84, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunckel:1901:I

[Gun01] John E. Gunckel. The index. Transactions of the American Fisheries Society, 30(1):62–65, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1900:HGC

[Hen00] James A. Henshall. Hints on grayling culture. Transactions of the American Fisheries Society, 29(1):109–117, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1901:PHF

[Hen01] James A. Henshall. Practical hints on fish culture. Transactions of the American Fisheries Society, 30(1):101–104, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1902:FGF

[Hen02] James A. Henshall. Food and game fishes of the Rocky Mountain region. Transactions of the American Fisheries Society, 31(1):74–88, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1903:BFF

[Hen03] James A. Henshall. Blood as fish food. Transactions of the American Fisheries Society, 32(1):63–70, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1904:EFF

[Hen04] James A. Henshall. Experiments in feeding fry. Transactions of the American Fisheries Society, 33(1):76–81, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1905:PFI

[Hen05] James A. Henshall. On the protection of fish in inland waters. Transactions of the American Fisheries Society, 34(1):139–147, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbard:1903:TGB

[Hub03] Waldo F. Hubbard. Transportation of green brook trout and salmon eggs, relative to the capacity of the two species of eggs to bear transportation or rough usage. Transactions of the American Fisheries Society, 32(1):79–89, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1900:CSF

[Jam00] B. W. James. The close season for fish. Transactions of the American Fisheries Society, 29(1):99–105, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1901:NPL

[Jam01] Bushrod Washington James. New Pennsylvania laws for the protection of food fish. Transactions of the American Fisheries Society, 30(1):144–151, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1904:RWD

[Joh04] F. M. Johnson. Resume of work done during the past year in the raising of western charr in Eastern waters. Transactions of the American Fisheries Society, 33(1):193–221, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1906:CFF

[Joh06] F. M. Johnson. The care and feeding of fry in pond life. Transactions of the American Fisheries Society, 35(1):230–232, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Joslyn:1905:PCC

[Jos05] C. D. Joslyn. The policy of ceding the control of the Great Lakes from State to National supervision. Transactions of the American Fisheries Society, 34(1):217–222, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lamkin:1900:SHL

[Lam00] J. Bayard Lamkin. The spawning habits of the large-mouth black bass in the South. Transactions of the American Fisheries Society, 29(1): 129–138, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lamkin:1902:FPB

[Lam02] J. B. Lamkin. A few points on the black bass — for discussion. Transactions of the American Fisheries Society, 31(1):147–152, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lambert:1904:PBP

[Lam04] E. M. Lambert. A plan for bass ponds. Transactions of the American Fisheries Society, 33(1):143–151, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1903:PLM

[Lea03] J. L. Leary. Propagation of large-mouth black bass at San Marcos Station. Transactions of the American Fisheries Society, 32(1):131– 191, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1904:CPP

[Lea04] J. L. Leary. Corstruction of ponds and pond cultural methods. Transactions of the American Fisheries Society, 33(1):139–142, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1906:QQB

[Lea06] John L. Leary. Questions for the question box, American Fisheries Society. Transactions of the American Fisheries Society, 35(1):225– 229, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1907:PFV

[Lea07] John L. Leary. Planting fish vs. fry. Transactions of the American Fisheries Society, 36(1):140–153, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1908:DSM

[Lea08] John L. Leary. Description of San Marcos Station with some of the methods of propogation in use at that station. Transactions of the American Fisheries Society, 37(1):75–81, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovejoy:1903:FFW

[Lov03] Samuel Lovejoy. Fish on the farm — what species to select. Transactions of the American Fisheries Society, 32(1):116–125, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovejoy:1904:WSB

[Lov04] Samuel Lovejoy. What I have seen of black bass. Transactions of the American Fisheries Society, 33(1):170–172, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1902:HCB

[Lyd02] Dwight Lydell. The habits and culture of the black bass. Transactions of the American Fisheries Society, 31(1):45–73, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1904:SNC

[Lyd04] Dwight Lydell. Some notes in connection with the bass work at Mill Creek Station. Transactions of the American Fisheries Society, 33(1): 152–169, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1906:BMC

[Lyd06a] Dwight Lydell. The bass at the Mill Creek Station. Transactions of the American Fisheries Society, 35(1):171–181, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1906:BPS

[Lyd06b] Dwight Lydell. The best paint for screens. Transactions of the American Fisheries Society, 35(1):187, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:1901:BTD

[Mar01] M. C. Marsh. The brook trout disease. Transactions of the American Fisheries Society, 30(1):66–81, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:1902:BTD

[Mar02] M. C. Marsh. The brook trout disease and cement ponds. Transactions of the American Fisheries Society, 31(1):107–115, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:1903:FAF

[Mar03] M. C. Marsh. A fatality among fishes in water containing an excess of dissolved air. Transactions of the American Fisheries Society, 32(1): 192–207, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:1904:DSC

[Mar04] M. C. Marsh. Danger in shipping cans: Notes of experiments made at the Cold Spring Harbor Hatchery of the New York Forest and Game Commission. Transactions of the American Fisheries Society, 33(1): 53–54, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mead:1900:BHG

[Mea00] A. D. Mead. The breeding habits and growth of the clam. Transactions of the American Fisheries Society, 29(1):176–179, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mead:1901:ELC

[Mea01] A. D. Mead. Experiments in lobster culture. Transactions of the American Fisheries Society, 30(1):94–100, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mead:1903:RAL

[Mea03] A. D. Mead. Recent advances in lobster culture. Transactions of the American Fisheries Society, 32(1):58–62, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mead:1905:PLC

[Mea05] A. D. Mead. The problem of lobster culture: Experimental ‘work of’ the Rhode Island Commissioners of Inland Fisheries. Transactions of the American Fisheries Society, 34(1):156–173, January 1905. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1904:YWF

[Mee04] W. E. Meehan. A year’s work of the fisheries interest in Pennsylvania. Transactions of the American Fisheries Society, 33(1):82–103, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1905:FC

[Mee05] W. E. Meehan. Frog culture. Transactions of the American Fisheries Society, 34(1):257–264, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1906:FDP

[Mee06a] W. E. Meehan. Fish distributed by Pennsylvania from January 1, 1906, to July 1, 1906. Transactions of the American Fisheries Society, 35(1):97, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1906:NTJ

[Mee06b] W. E. Meehan. New type of jars for hatching. Transactions of the American Fisheries Society, 35(1):169–170, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1907:SWD

[Mee07] W. E. Meehan. The shad work on the Delaware River in 1907 and its lessons. Transactions of the American Fisheries Society, 36(1): 105–118, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1904:PES

[Moo04] H. F. Moore. Progress of experiments in sponge culture. Transactions of the American Fisheries Society, 33(1):231–243, January 1904. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nevin:1901:PMW

[Nev01] James Nevin. The propagation of muskellunge in Wisconsin. Transactions of the American Fisheries Society, 30(1):90–93, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:1907:NIF

[Ols07] O. T. Olsen. The need of an international fisheries society. Transactions of the American Fisheries Society, 36(1):252–254, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OMalley:1905:SSA

[O’M05] Henry O’Malley. Salt solution as an aid to fish culture. Transactions of the American Fisheries Society, 34(1):49–56, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1901:MCF

[Par01] J. C. Parker. Man as a controlling factor in aquatic life. Transactions of the American Fisheries Society, 30(1):48–61, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peabody:1904:THB

[Pea04] George F. Peabody. A tribute to Hon. E. E. Bryant. Transactions of the American Fisheries Society, 33(1):43–52, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pond:1902:RLA

[Pon02] Raymond H. Pond. The role of the larger aquatic plants in the biology of fresh water. Transactions of the American Fisheries Society, 31(1): 89–96, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reighard:1906:ILP

[Rei06] Jacob Reighard. On the identification for legal purposes of mutilated or dressed specimens of whitefish and herring from the Great Lakes. Transactions of the American Fisheries Society, 35(1):47–58, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ripple:1908:ERS

[Rip08] Robert Ripple. Experiments in rearing small-mouth black bass. Transactions of the American Fisheries Society, 37(1):126–131, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosenberg:1906:SEP

[Ros06] Albert Rosenberg. Some experiments in the propagation of rainbow trout. Transactions of the American Fisheries Society, 35(1):87–90, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roth:1906:FR

[Rot06] Filibert Roth. The fisherman and reforestation. Transactions of the American Fisheries Society, 35(1):164–168, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raveret-Wattel:1907:ORV

[RW07] E. Raveret-Wattel. Observations regarding the varieties of rainbow trout and the time of the spawning of this species. Transactions of the American Fisheries Society, 36(1):119–127, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seagle:1902:SRR

[Sea02] George A. Seagle. Some remarks on the rainbow trout, the time for planting, etc. Transactions of the American Fisheries Society, 31(1): 138–146, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1904:JPF

[Smi04] Hugh M. Smith. Japan, the paramount fishing nation. Transactions of the American Fisheries Society, 33(1):111–138, January 1904. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1905:RSC

[Smi05a] H. M. Smith. Remarks on sponge cultivation. Transactions of the American Fisheries Society, 34(1):256, January 1905. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1905:ICF

[Smi05b] Hugh M. Smith. The International Congress of Fisheries at Vienna, 1905. Transactions of the American Fisheries Society, 34(1):199–200, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1906:FLB

[Smi06] Hugh M. Smith. Fishery legislation before the Fifty-Ninth Congress, First Session. Transactions of the American Fisheries Society, 35(1): 91–96, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1907:SOE

[Smi07] Hugh M. Smith. Some observations on European fisheries and fish culture. Transactions of the American Fisheries Society, 36(1):170– 183, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1900:SHL

[Sto00] Livingston Stone. The spawning habits of the lake sturgeon: (Acipenser rubicundus). Transactions of the American Fisheries Society, 29(1):118–128, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1901:SHL

[Sto01] Livingston Stone. Sturgeon hatching in the Lake Champlain Basin. Transactions of the American Fisheries Society, 30(1):137–143, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranahan:1900:LFV

[Str00] J. J. Stranahan. Lack of fertilization vs. arrested segmentation. Transactions of the American Fisheries Society, 29(1):173–175, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranahan:1902:FCF

[Str02] J. J. Stranahan. Fish culture on the farm. Transactions of the American Fisheries Society, 31(1):130–137, January 1902. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranahan:1903:TMP

[Str03] J. J. Stranahan. Three main points necessary to successful bass culture. Transactions of the American Fisheries Society, 32(1):126–130, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranahan:1906:ABB

[Str06] J. J. Stranahan. Assorting brood black bass to prevent cannibalism. Transactions of the American Fisheries Society, 35(1):183–186, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranahan:1907:TVP

[Str07] J. J. Stranahan. Theory vs. practical tests. Transactions of the American Fisheries Society, 36(1):246–247, January 1907. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranahan:1908:SPS

[Str08] J. J. Stranahan. Some peculiarities in spawning habits of large-mouth black bass. Transactions of the American Fisheries Society, 37(1): 157–159, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sykes:1902:IPR

[Syk02] Arthur Sykes. Inbreeding pond-reared trout. Transactions of the American Fisheries Society, 31(1):116–124, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talbott:1905:PB

[Tal05] Henry Talbott. Potomac bass. Transactions of the American Fisheries Society, 34(1):124–132, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1900:BTF

[Tho00] W. T. Thompson. Brook trout fry; a resume of methods. Transactions of the American Fisheries Society, 29(1):139–153, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1901:BTN

[Tho01] W. T. Thompson. Brook trout notes. Transactions of the American Fisheries Society, 30(1):152–158, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1902:FEG

[Tho02] W. T. Thompson. Feeding: Its effect on growth and egg production. Transactions of the American Fisheries Society, 31(1):125–129, January 1902. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1903:GT

[Tho03] W. T. Thompson. The golden trout. Transactions of the American Fisheries Society, 32(1):208–217, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1904:FRS

[Tit04] John W. Titcomb. Fishery reminiscences in South America. Transactions of the American Fisheries Society, 33(1):244, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1905:RFS

[Tit05a] J. W. Titcomb. Reminiscences of the fisheries in South America. Transactions of the American Fisheries Society, 34(1):117, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1905:PEF

[Tit05b] John W. Titcomb. Progress and experiments in fish culture during the past year in the Bureau of Fisheries. Transactions of the American Fisheries Society, 34(1):57–74, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1906:PEF

[Tit06] John W. Titcomb. Progress and experiments in fish culture in the Bureau of Fisheries during the Fiscal Year 1906. Transactions of the American Fisheries Society, 35(1):98–122, January 1906. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Townsend:1907:CFS

[Tow07] Charles H. Townsend. The cultivation of fishes in small ponds. Transactions of the American Fisheries Society, 36(1):128–139, January 1907. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1908:TAT

[Wag08] George Wagner. Tullibee (Argyrosomus tullibee Richardson) as a fish of economic importance. Transactions of the American Fisheries Society, 37(1):122–125, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1903:SNF

[War03] H. B. Ward. Some notes on fish food in the lakes of the Sierras. Transactions of the American Fisheries Society, 32(1):218–220, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1908:SPM

[War08] Henry B. Ward. Some points in the migration of Pacific salmon as shown by its parasites. Transactions of the American Fisheries Society, 37(1):92–114, January 1908. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whish:1903:CVS

[Whi03] John D. Whish. Commercial values: Some notes of a study of the work of the Forest, Fish and Game, Commission of New York State. Transactions of the American Fisheries Society, 32(1):112–115, January 1903. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whish:1905:PNB

[Whi05] John D. Whish. The passing of the native brook trout. Transactions of the American Fisheries Society, 34(1):90–115, January 1905. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wood:1900:EOC

[Woo00] C. C. Wood. Exhibition of original contrivances for use at brook trout hatcheries; with description and discussion. Transactions of the American Fisheries Society, 29(1):51–68, January 1900. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wood:1901:QWF

[Woo01] C. C. Wood. The quality of the water a factor in rearing trout fry. Transactions of the American Fisheries Society, 30(1):105–113, January 1901. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Worth:1904:RHS

[Wor04] S. G. Worth. The recent hatching of striped bass and possibilities with other commercial species. Transactions of the American Fisheries Society, 33(1):223–230, January 1904. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).